OpenStudy (anonymous):

You make $10.00/hour and work 40 hours a week. You invest $4000 dollars at an interest rate of 5% APR for a year. How much interest did that savings earn in a year in terms of hours of work?

OpenStudy (tkhunny):

4000*0.05 = ??

OpenStudy (anonymous):

200

OpenStudy (tkhunny):

200/10 = ??

OpenStudy (anonymous):

20..

OpenStudy (tkhunny):

You're done.

OpenStudy (anonymous):

Lol. Thanks hun<3

OpenStudy (tkhunny):

It might make more sense with the units. \($4,000⋅0.05=$200\) \($200\dfrac{1\;hr}{$10.00}=20\;hr\)